# Examining Inequality in Park Quality for Promoting Health Across 35 Global Cities

### v1

@@ 2.2 Park Health Scores
Figure 3: Park spaces in an average 8 ha park dedicated to nature-appreciation & physical activities in North America and Europe. North American cities offer more spaces for physical activities, whereas European cities devote more space to nature-appreciation. Standard box plot with Q1, median, Q3; whiskers at 1.5×1.5$\times$1.5 × interquartile range.
European parks offer most spaces for nature appreciation, and North-American for physical activities. The most notable difference revealed by the average park offering analysis is between European and North American cities in the physical and nature-appreciation categories, as shown in Figure 3. European urban parks are primarily dedicated to nature appreciation, while North American parks allocate a larger proportion of spaces to physical activities. This highlights the priorities in urban planning across regions and reveals the different roles parks play in various cultures.
2.3 Inequality of Park Scores Within and Across Cities
Centralization of Park Quality Within Cities
Turning our attention to the locations of parks that are great for doing health-promoting activities, we found that parks in the city center tend to score higher not only in cultural and social activities but also in physical and nature-appreciation activities. Specifically, we categorized all parks into four quartiles based on their proximity to the city center, from the innermost 25% (Q1) to the outermost 25% (Q4) (Figure 4). This analysis revealed a clear trend: the closer a park is to the city center, the higher its score in health-promoting activities.
Interestingly, the most significant discrepancies were observed in nature-appreciation scores, challenging the assumption that suburban areas, due to less crowding, would have more natural spaces (Table 5). This phenomenon, described as the “Savannah Trap” by Montgomery [153], highlights a common oversight in suburban planning where urban planners regularly fail to create biodiverse spaces in the suburbs, instead inadvertently creating monotonous, savannah-like expanses between residential blocks. Our findings, detailed further in Section A.3 in the Appendix, were consistent across cities on every continent, with only two exceptions in South America (Buenos Aires and Rio de Janeiro, where the urban topology is influenced by the sea), demonstrating the widespread applicability of these trends.
Figure 4: Park health scores by quartiles of the distance to the city center. Q1 are the inner city parks, and Q4 are the parks that are most distant from the city center. The scores decrease monotonously with each quartile further away from the city center, meaning that the farther a park is to the city center, the lower is its score in health-promoting activities. The effect is strongest for nature-appreciation and is least prominent for environmental activities. Standard box plot with Q1, median, Q3; whiskers at 1.5×1.5$\times$1.5 × interquartile range.
Table 5: Park health scores by quartiles determined using the distance to the city center. Q1 are the inner city parks, and Q4 are the parks that are most distant to the city center. On the left, we show the mean health scores for each quartile; on the right, we show the $p$-values for each null hypothesis $H_{0}:Q_{i}=Q_{j}$ stating that the mean scores for two subsequent quartiles $i$ and $j$ are equal. All null hypotheses can be refuted with high significance levels ($p<0.001,\ast\ast\ast$), with the exception of the difference between Q3 and Q4 in the environmental category, where the significance is $p=0.003,\ast\ast$.
Mean Health Scores $p$-values of $H_{0}$
Activity Category Q1 Q2 Q3 Q4 Q1 = Q2 Q2 = Q3 Q3 = Q4
Physical 0.382 0.297 0.160 0.068 0.000 0.000 0.000
Nature-appreciation 0.107 -0.211 -0.497 -0.580 0.000 0.000 0.000
Environmental 0.020 -0.076 -0.156 -0.186 0.000 0.000 0.003
Social 0.289 0.137 -0.004 -0.088 0.000 0.000 0.000
Cultural 0.260 0.096 0.010 -0.022 0.000 0.000 0.001
Disparities in the Offered Activities Across Cities
Following our examination of how park scores differ within cities, we quantified the disparities in park health scores across different cities. To that end, we used an inequality metric (refer to Equation 3 in Methods Section 4.6), which ranges from 00 (indicating uniform distribution of facilities) to $1$ (indicating absolute disparity).
Figure 5: Disparities in Park Health Scores by activity categories (left) and continents (right). While the disparities of the cities converge to a relatively uniform level between $0.15$ and $0.19$ across all continents, different activities exhibit varying levels of disparities. Physical facilities have the most uniform provision, whereas social facilities show the largest disparities. Standard box plot with Q1, median, Q3; whiskers at 1.5×1.5$\times$1.5 × interquartile range. The white cross indicates the mean value.
The box plots in Figure 5 aggregate the disparity scores by continent and activity category. The individual values are tabulated in Table 6. Generally, we found low disparities, with an average score of $0.165$ across all categories and cities. Disparities were notably low in physical activities ($\mu=0.135,\sigma=0.039$), suggesting a consistent distribution of physical activity facilities across parks. Conversely, Social activities exhibited the highest disparity ($\mu=0.222,\sigma=0.073$), indicating significant variation in social amenities among parks within cities. Nature-appreciation activities displayed a slightly higher average disparity ($\mu=0.157$) than physical activities but with the least variation ($\sigma=0.032$), indicating a relatively consistent focus on nature-appreciation facilities across the cities studied.
We hypothesize that equitable distribution of amenities and facilities supporting different activity categories is generally beneficial from a societal fairness perspective. Interestingly, the disparities of the cities converge to a relatively uniform level between $0.15$ and $0.19$ across all continents (Figure 5b). This indicates that there is no clear divide between the different regions of the world when it comes to fairness aspects of park quality. We use Table 6 to investigate this phenomenon on a city level.
@@ 2.3 Inequality of Park Scores Within and Across Cities
Table 6: Comparison of disparity scores across cities by activity category. The mean value over all activity categories determines the table order. Lower values signify a more uniform offering of activity type across a city, whereas higher values indicate a greater disparity in that offering. Values lower or higher than $1$ standard deviation from the mean value a marked in gray and bold, respectively.
@@
Cities Providing a Range of Activities. Tokyo, Paris, Auckland, Buenos Aires, and Hong Kong were found to have the most equitable provision of park offerings across our five health categories. Tokyo and Paris stood out, with five and four categories, respectively, being below one standard deviation from the mean disparity. These minimal disparities in park facilities may result from centralized park management and legislation promoting community involvement in Paris[154], and equitable public service provision in Tokyo[155].
Cities Focused on Specific Activities. At the other end of the spectrum in Table 6, Moscow, St. Petersburg, Stockholm, Rio de Janeiro, and Copenhagen exhibit pronounced disparities in park facilities. Stockholm has the highest individual disparity score in the cultural dimension ($0.514$), indicating a considerable centralization of cultural facilities in only a few parks in the Swedish capital. Copenhagen’s inequality in park facilities is especially high in the cultural and social dimensions, particularly in central districts, raising concerns that the city’s transformation in recent decades has overly catered to the affluent[156]. In Rio de Janeiro, disparities were observed across social, environmental, and cultural activities, aligning with existing research on health inequalities[157] and reflecting the city’s deep-rooted socioeconomic segregation[158].
@@ 3 Discussion
Urban parks offer a variety of health benefits that extend beyond the typical physical and mental advantages usually highlighted. By discriminating six types of activities people do in parks, we created a detailed taxonomy of park activities and their associated health benefits as found in the medical literature. Our efforts to provide a balanced review unveiled a clear bias towards physical and nature-appreciation activities in the literature. We used this taxonomy to evaluate parks globally based on their support for five of the six activity categories (data to evaluate the mind-body category was unavailable on OSM). By providing individual scores for each activity, we were able to give useful insights about the parks. This allows citizens to identify parks that best suit their preferences and needs, while city authorities can determine which parts of the city may be lacking in parks with specific facilities.
Despite all the positive aspects of parks, disparities in access to urban greens is a prevalent theme in academic literature[159, 160, 161, 162, 5, 163]. Our city-level results also reveal disparities in park facilities, with a clear geographic trend: parks located in the heart of the city are better equipped for health-promoting activities than those on the outskirts. This pattern holds true across all types of activities, even ones that intuitively would be more available in less populated areas (i.e., outskirts), like nature appreciation or environmental activities. This finding aligns with the work of other researchers, such as Wolch et al., who conclude (mostly based on Western studies) that the distribution of parks in a city disproportionately benefits affluent communities[159]. Furthermore, it has been shown that the urbanization of the past two decades has resulted in considerable greening in the downtown areas of cities, whereas the suburbs have not seen the same level of investment [161]. Our results show, additionally to the diminished access to parks in the suburbs described in the literature[161, 159], that the quality of these parks is lacking compared to the city center parks. This is a call to action for urban planners to shift their focus to improving these suburban parks. Remarkably, the greatest contributions could be made by creating more natural spaces that support nature-appreciation activities, followed by social activities.
We also investigated overall disparities of facilities across cities. Here, the findings are generally positive in the sense that the overall disparity scores were relatively low and did not display a pronounced geographic variability on a global scale as in previous studies[162, 164, 165]. This finding can be summarized in a way that while absolute access to greenery has been shown to have quite a high level of inequality, the relative inequality regarding access to health-promoting activities in parks is less severe. Interestingly, the five cities showing the greatest disparities in offerings of activity types are spread across four different continents, indicating that overall disparity in the quality of parks is a prevalent issue worldwide. This finding suggests that just urban planning is not a matter of financial resources, but a question of municipal decision-making and community involvement[154, 166, 167]. A positive trend can be observed in Asia and Oceania, which have the most equitable provision of park facilities. We found that physical activities were the most evenly distributed, indicating that they receive more attention from urban planners [25, 28]. This is in line with the focus of medical literature on the health benefits of physical activity in our literature review (cf. Section 4.3). On the other hand, dedicated social activities showed the highest inequality, indicating a clear need for improvement in these areas in the respective cities[168]. We hope that, based on these insights, urban planners will recognize the need to look beyond physical activities and put more emphasis on improving parks in other ways so to benefit all the citizens (e.g., elderly, who might not be able to engage in physical activities).
While creating high-quality urban green areas might be a balancing act with potential negative effects such as gentrification arising[159], our research can serve as a roadmap for urban planners about which parks to improve with which kind of specific facilities. The ongoing urbanization will not provide them with an easy task, and the obesity crisis coupled with increasing costs of living in cities puts planners into a dilemma: should the available space be used for housing, parks, or other services[169]? Irrespective of such considerations, improving the existing parks with more facilities for health-promoting activities will be foundational for public health[3, 22], especially in the outskirts of the cities. This needs to be accompanied by efforts to make parks and their facilities accessible for everyone[170].
In the broader context of urban health, our research adds to earlier studies that explored the environmental benefits of urban green spaces, such as cooling effects[171, 172] and reduced air pollution[173, 174]. Our focus on activities provides an opportunity to merge these findings, for example, in the form of recommendations for the best times to enjoy activities in certain parks when the environmental conditions are adequate to do so.
Limitations and Future Directions
@@
The first limitation of our work was that activities not requiring specialized facilities could not be accurately considered using our method. While there are approaches to understanding the relationship between open spaces in parks and their usage[175], we have refrained from speculating about such usage to avoid any cultural bias in our park evaluations. This meant we had to remove mind-body activities like yoga from our main analysis, as they often do not need designated areas in city parks. Similarly, our data does not typically account for temporary cultural events like music festivals held in parks.
The quality of parks was measured using map data from OSM, under the assumption that if facilities and spaces that support certain activities are present, they will be utilized for those purposes. However, as acknowledged in the previous limitation, not all activities are directly linked to specific facilities, and the intensity of use cannot be determined using OSM cartographic data alone. This gives rise to instantiating our activity taxonomy (Table 1) using a different data source, which could more accurately reflect actual usage. We demonstrated that the OSM park scores are closely linked to activities captured in Flickr photos, but other data could also be used to approximate park usage. For instance, detailed mobile traffic data could offer intriguing insights into how parks are used[176, 43, 177].
Our scores for each city’s parks are adjusted to reflect the unique circumstances of that city, making direct comparisons between different cities’ park scores impermissible. This property prevents some comparative analyses but also avoids comparing vastly different situations as each city has its own cultural, historical, climate, and geographical factors that promote or hinder having high-quality parks. By individually scoring each city, we tailor our ratings to the specific needs of each city, helping urban planners address the most urgent improvements in their city.
We focused on urban parks, excluding other urban green spaces like gardens, street trees, and green roofs, which also promote urban health. The choice of parks was because they offer a wider range of activities compared to the more specialized functions of, e.g., gardens—primarily dedicated to the appreciation of nature—and street trees, which enhance walkability. Parks indeed have a premier role in urban life due to their public accessibility[178], and the commission of their facilities usually falls under the influence of the city administration, meaning our findings about any shortcomings could lead to actionable improvements.
@@
By analyzing snapshots of the OSM database, we captured the state of park amenities and facilities in fall 2023. Like all urban spaces, parks are in constant flux, which raises the question of how parks have evolved over time[5]. This is particularly relevant in the context of the COVID-19 pandemic, when many cities had to reconsider their public health strategies, possibly leading to significant changes in park facilities[163, 179]. By examining historical data from OSM, we could use our methodology to track these changes in park facilities. However, a key challenge would be differentiating between actual changes in park development and potential temporal variations in OSM data quality.
Our study primarily examined park facilities that encourage health-promoting activities. However, we did not investigate how these facilities are actually utilized. Therefore, a key area for future research could be analyzing how park quality directly affects people’s health. Although it would be unrealistic to gather medical data worldwide to determine the health impact of individual parks, more targeted studies might be possible in certain areas. For example, prescription data that are available for some countries, such as England, could be used to estimate various medical conditions[180, 181]. It would also be worthwhile to further investigate the causal relationships between the temporal changes of parks resulting in better health (e.g., fewer prescriptions).

### v2

@@ 2.2 Scoring Parks for Health Support
Bottom: Average space in 8-hectare parks dedicated to nature and physical activities in North America and Europe. North American parks support physical activity more, while European parks provide more space for nature appreciation. Standard box plot with Q1, median, Q3; whiskers at $1.5\times$ interquartile range.
2.3 Cities Differ in Health-Related Activity Support
Our analysis revealed systematic variations in how urban parks support health-promoting activities, with differences occurring both between cities and within individual urban areas. These variations reflect distinct urban planning priorities, geographic influences, and equity considerations that shape park design and resource allocation.
Estimating Park Offerings Through 8-Hectare Model Parks
To compare park features systematically across cities, we constructed statistical models of typical 8-hectare parks: a size representing neighborhood parks and matching our dataset’s average. Figure 2 (top) shows the expected number of elements and area dedicated to each activity type in a statistically average 8-hectare park for each city. We computed these values using parameters from linear regression models developed for each city (Equation 1). While we fitted separate models per city, making raw health scores not directly comparable between cities, the models enabled comparison of hypothetical average parks across urban contexts.
Regional Priorities: European Nature Focus vs. North American Physical Activity
The most notable regional contrast lies in emphasis on physical activity versus nature appreciation, as shown in Figure 2 (bottom). Parks in European cities tended to provide more space for nature appreciation, while North American cities allocated more area to physical activities. These differences reflect broader urban planning patterns and suggest that parks serve different roles depending on regional goals and cultural expectations.
Cultural Features: Universally Limited Support
Across all regions, cultural features appeared least frequently in parks. In terms of elements, nature appreciation appeared most often, followed by social and environmental activities. Cultural features were the least common. This pattern suggests that while nature and movement receive widespread support, cultural health-promoting activities are deprioritized globally in urban park design.
2.4 Park Offerings are Unequally Distributed Within Cities
Figure 3: (a) Park scores by distance quartile. Q1 includes parks closest to the city center; Q4 includes those farthest away. Scores decline steadily with distance, especially for nature appreciation. Environmental features show the least decline. Standard box plot with Q1, median, Q3; whiskers at $1.5\times$ interquartile range. (b, c) Inequality in park health scores. Disparities across activities and continents are shown. While total disparity ranges are similar across regions, some activities, such as social and cultural, show much higher inequality than others.
We examined how park scores vary within cities by grouping parks into quartiles based on distance from city centers, from the innermost 25% (Q1) to the outermost 25% (Q4). This approach allowed us to systematically assess whether proximity to urban centers influences park offerings.
Inner-City Parks Consistently Outperform Peripheral Ones
Across all activity types, parks in central areas scored higher. Figure 3a demonstrates that park scores dropped steadily with each quartile further from the center. This pattern held across most cities on all continents, with exceptions in Buenos Aires and Rio de Janeiro, where geography and coastal form influence green space layout.
Nature Appreciation Shows Strongest Distance-Related Decline
The trend was most pronounced for nature appreciation. This finding challenges assumptions that outer areas naturally support greener, more restorative environments and echo Montgomery’s concept of the “Savannah Trap[35]”, where suburban areas include empty, open land that does not support social or ecological use.
2.5 Cities Vary in Ensuring Equitable Park Access
@@
Table 2: The disparity scores by city and activity type are generally low, but vary based on the activity category. Cities are ordered by their overall average disparity. Lower values indicate a more even distribution of features. Scores one standard deviation below or above the mean are marked in gray and bold, respectively.
@@
We assessed equity in park offerings within cities using a disparity index (Equation 3) ranging from 0 (complete equality) to 1 (maximum inequality). Box plots in Figure 3b–c show average disparities by activity and region, with detailed values in Table 2.
Activity-Specific Disparity Patterns
Disparities varied by activity type: physical activities showed the most even distribution ($\mu=0.135,\sigma=0.039$), while social activities exhibited the highest disparity ($\mu=0.222,\sigma=0.073$), indicating that some parks offer rich social features while others lack them entirely. Nature appreciation had moderate average disparity ($\mu=0.157$) but low variation ($\sigma=0.032$), suggesting consistent global trends. Overall disparities averaged $0.165$, and were similar across continents (0.150.150.190.19), suggesting that inequality in park design is not limited to one region.
Cities Achieving Balanced Activity Support
Tokyo, Paris, Auckland, Buenos Aires, and Hong Kong demonstrated the most equitable park offerings across activities. Tokyo and Paris particularly excelled, with four or more activity categories showing disparities below one standard deviation from the mean. These results may reflect centralized park governance, strong public policy, and active community involvement in cities such as Paris [36] and Tokyo [37].
Cities with Concentrated Park Resources
In contrast, Moscow, St. Petersburg, Stockholm, Rio de Janeiro, and Copenhagen showed high disparities in park offerings across categories. Stockholm exhibited the highest single-category disparity in cultural features ($0.514$), suggesting cultural resources were concentrated in a few parks. Copenhagen showed high disparity in both cultural and social categories, potentially reflecting gentrification[38]. Rio de Janeiro displayed widespread disparities across social, environmental, and cultural categories, aligning with prior findings on spatial health inequality[39].
These findings demonstrate that while park inequality exists globally, specific cities have successfully implemented more equitable approaches to supporting diverse health-promoting activities through their urban park systems.
@@ 3 Discussion
Urban parks offer health benefits that go beyond the usual physical and mental wellbeing. However, our global analysis reveals that urban parks require strategic improvement focused on geographic equity and activity diversity to maximize their health benefits.
3.1 Main Findings
By identifying six types of activities people do in parks, we created a detailed taxonomy of park activities and used this to evaluate parks worldwide based on their support for five of the six activity categories (data to evaluate the mindfulness category was not available on OSM). By giving individual scores for each activity, we identified three critical areas that demand attention from landscape architects, urban designers, and policymakers.
Geographic Inequities Demand Immediate Action
Our findings show a clear and concerning geographic trend: parks in city centers are better equipped for health-promoting activities than those on the outskirts. This pattern holds across all types of activities, including those that might be expected to be more common in less populated areas, such as nature appreciation or environmental activities. This geographic disparity contradicts basic expectations and creates significant barriers to equitable health access.
This finding aligns with previous research. Wolch et al. found that park distribution often favors more wealthy communities[31]. Other studies have shown that recent urban development policies have led to significant greening in city centers, while suburban areas have received less investment[40]. Our results expand on these works by showing not only reduced access to parks in suburban areas[40, 31], but also a lack of amenities and spaces for health-promoting activities.
This evidence creates a clear call to action for urban planners to shift their focus to improving suburban parks, as well as parks outside urban centers, where significant and often overlooked recreation gaps are often found [41]. However, creating high-quality urban green spaces can be a balancing act, with potential downsides such as gentrification[34, 31]. Designers must think carefully about the urban context when improving parks in ways that may change their role[34].
Activity-Type Inequities Show Mixed but Encouraging Patterns
We also examined disparities in the health potential of parks among cities by activity type. The findings are generally encouraging: overall disparity scores were low and did not show pronounced global variation, in contrast to earlier studies[42, 43]. While previous work has shown high inequality in absolute access to greenery, our results suggest that relative inequality in access to health-promoting park activities is less severe.
The distribution patterns vary significantly by activity type, revealing both strengths and critical gaps. Physical activity infrastructure was the most evenly distributed, likely reflecting its prominence in public health discourse and urban policy[18]. This aligns with the medical literature’s emphasis on physical activity benefits (Table 7). In contrast, facilities for social activities were the most unequally distributed, revealing a key area for improvement[44].
The global nature of this challenge is evident from our data: the five cities with the greatest disparities in park offerings span four continents, indicating that uneven provision of well-equipped parks is not a concentrated issue. This suggests that just urban planning is not only a matter of financial resources but also one of municipal priorities and community engagement[36, 45]. A positive trend can be observed in the Asian countries of our study and Oceania, which have the most fair provision of park facilities.
Our Findings Provide Actionable Insights for the Design of Well-equipped Parks
To understand the relevance of our work for urban planning, we conducted 30-minute semi-structured interviews with three domain experts (see Section B.8). We reached out to practitioners with 12–30 years of professional experience from diverse backgrounds: a lecturer and urban designer (E1, male, Hong Kong), an urban designer and master planner (E2, female, North America), and a municipal park development manager (E3, male, United Kingdom). The experts confirmed that our quantitative approach offers significant practical value for urban design decisions. They identified several ways our findings could support urban design: reducing reliance on subjective park assessments (E1), serving as a quantitative baseline for master planning (E1 and E2), and providing an evidence-based alternative to institutional knowledge and practice-specific heuristics (E1). The urban designers emphasized that the visualization of park scores is already useful to “identify the gaps regarding what is present in terms of offerings in a district (E1)”, especially when aiming to ensure fair access to parks. This was echoed by the park manager, who stated they would use the park scoring to guide development priorities, and as supporting evidence to gain leadership backing or external funding. When shown the taxonomy of activities, they noted that most of their projects addressed physical and nature-appreciation activities, and “[the activity taxonomy] helps us think about what it is we’re doing and how and how some categories have been neglected (E3).”
The experts also highlighted the importance of the way our scores are normalized to the specific context of each city. Our park scores are normalized to reflect the unique context of each city, making direct comparisons between cities hard. While this limits inter-city comparisons, it avoids drawing misleading conclusions across culturally, geographically, and climatically diverse settings. For example, E1 cited culturally different approaches to park provision in the form of “three-dimensional parks” in high-density environments like Hong Kong, while E2 brought up the challenges in adopting urban transformation manuals from Western cities in regions with extreme climates such as the Middle East.
3.2 Limitations and Future Directions
Despite the well-known benefits of parks, lack of fair access to urban green spaces remains a common theme in academic literature[31, 40, 42, 2, 46]. We focused only on urban parks, excluding other types of urban green spaces such as gardens, street trees, and green roofs, which also contribute to urban health. Parks were selected because they support a wider range of activities than more specialized spaces, are generally publicly accessible, and are typically maintained by municipal administrations, making any identified shortcomings more actionable for policy and planning.
@@ 3.2 Limitations and Future Directions
Relying on map data limits our analysis to activities that can be linked to specific elements and spaces within parks. While some studies explored the relationship between open spaces and their use[47], we avoided making assumptions about the usage of informal spaces to reduce the risk of cultural bias in our park evaluations. As a result, we excluded mindfulness activities from our main analysis, as they often do not require designated areas in city parks. Likewise, our data does not account for temporary cultural events, such as music festivals held in parks.
Park offerings were scored under the assumption that the presence of facilities and spaces enabling certain activities is a necessary condition for realizing specific health benefits. However, their presence alone may not be enough to deliver these benefits in practice. Furthermore, not all activities are directly tied to specific infrastructure, and OSM cartographic data alone cannot capture the intensity of use. We did not distinguish between the quality of maintenance or design of individual elements, such as benches being equipped with backrests, or the size, layout and species of trees and other plants.
@@
Like all urban spaces, parks are in constant flux, which raises the question of how parks have evolved over time [2]. This became clear during the COVID-19 pandemic, when many cities had to rethink their strategies, possibly leading to big changes in park facilities with major social justice effects [48]. Looking at historical data from OSM is challenging, as it is difficult to distinguish between actual changes in park facilities and the increasing completeness of the OSM database.
A key direction for future research is to examine how park offerings directly influence health outcomes. While collecting global medical data to evaluate the health impact of individual parks is impractical, more targeted investigations may be feasible in specific contexts. For instance, prescription data available in some countries could be used to estimate the prevalence of certain medical conditions[49], and to explore potential causal links between changes in park infrastructure and improvements in public health over time.
We showed that the OSM park scores are closely linked to activities captured in Flickr photos, but other data could also be used to approximate park usage. For instance, detailed mobile traffic data could offer interesting insights into how parks are used[50, 30, 51]. This could also help reduce biases inherent in social media data sources like Flickr and OSM, which rely on actively submitted contributions from tech-savvy users, and sometimes out-of-city visitors rather than passive observations of citizen behavior.
Finally, by focusing on health-promoting activities, our work introduces avenues for health-related behavioral change[52, 20]. For instance, our activity-based approach could support context-aware recommendations[53], such as identifying the most suitable times to engage in specific park activities when conditions like temperature, air quality, or crowdedness are most favorable.